# Towards Cross-Tokenizer Distillation: the Universal Logit Distillation Loss for LLMs

### v2

@@ Abstract
Deploying large language models (LLMs) of several billion parameters can be impractical in most industrial use cases due to constraints such as cost, latency limitations, and hardware accessibility. Knowledge distillation (KD) offers a solution by compressing knowledge from resource-intensive large models to smaller ones. Various strategies exist, some relying on the text generated by the teacher model and optionally utilizing his logits to enhance learning. However, these methods based on logits often require both teacher and student models to share the same tokenizer, limiting their applicability across different LLM families. In this paper, we introduce Universal Logit Distillation (ULD) loss, grounded in optimal transport, to address this limitation. Our experimental results demonstrate the effectiveness of ULD loss in enabling distillation across models with different architectures and tokenizers, paving the way to a more widespread use of distillation techniques.
@@ 1 Introduction
A noticeable trend has emerged in NLP with the prevalence of large language models (LLMs) such as LLama Touvron et al. (2023a), Mistral Jiang et al. (2023), Falcon Almazrouei et al. (2023), GPT-NeoX Black et al. (2022), or Mixtral Jiang et al. (2024). Despite the high performance of LLMs Bubeck et al. (2023), the challenges associated with resource consumption and deployment complexity have become increasingly prominent due to hardware availability, cost, and latency bottlenecks. Several methods, including efficient decoding Leviathan et al. (2023); Ye et al. (2023), model recycling Lester et al. (2022), and size reduction Dettmers et al. (2023); Ma et al. (2023), have been proposed to address the need for faster and more efficient deployment of these models. In this paper, we concentrate on knowledge distillation (KD) Buciluundefined et al. (2006); Hinton et al. (2015), a widely adopted technique Sanh et al. (2020); Jiao et al. (2020); Mohammadshahi et al. (2022); He et al. (2023); Raman et al. (2023); Dasgupta et al. (2023), utilized by practitioners to distill the expertise of a larger teacher model into a compact student model, preserving maximum performance while reducing latency and memory footprint.
Over the past years, NLP researchers have extensively explored and applied knowledge distillation mostly on smaller student models derived from BERT Sanh et al. (2020); Jiao et al. (2020); Sun et al. (2020). These smaller student models maintain a similar architecture to the teacher, mirroring some of its blocks, hidden sizes, or relying on the same tokenizer. Two approaches can generally be considered: the white box approach, where researchers propose loss functions that directly compute similarities across layers, and the black box approach, offering flexibility, agnostic to teacher latent states. This black box approach can be readily implemented by practitioners through libraries and APIs, simplifying its adoption.
Figure 1: Ratio of student vocabulary contained in teacher tokenizer. (e.g., Bloomz tokenizer has 30.03% of Mistral’s vocabulary). For student and teacher model information see Sec. 4.2
@@
However, KD for generative models, those relying on encoder-decoder or decoder architectures, has received less attention. Recent research, predominantly focuses on synthetic data fine-tuning He et al. (2022); Kramchaninova and Defauw (2022); Ouyang et al. (2022), with less emphasis on refining loss functions in the black box approach. Thus far, the primary method for KD involves using text generated by the teacher model He et al. (2023); Hsieh et al. (2023), and optionally, augmenting it with logits, employing Kullback–Leibler divergence between the teacher logits distribution and student ones Mohammadshahi et al. (2022); Raman et al. (2023); Wang et al. (2020a). Although the logit distillation method yields significant improvements, it has been underutilized in recent studies due to its requirement for the student model to share the same tokenizer as the teacher (Sec. 2.4 and Fig. 2). Indeed, even when encoder-decoder or decoder-only models of varying sizes are available, they often do not share the same architecture nor the same tokenizer (Fig. 1), rendering Logit Distillation loss inapplicable. This raises the following research question:
How can we craft a Universal knowledge distillation loss within the black box approach, that is inherently versatile to teacher/student architectures and agnostic on tokenizers?
@@
    A universal logit distillation loss. We introduce a new loss, Universal Logit Distillation Loss (ULD loss), with virtually no assumptions on the teacher and student architectures. Our approach harnesses a closed-form solution of optimal transport, rendering it ideal for large-scale fine-tuning due to its fast computation.
@@
    Experimental Results. To showcase the consistent effectiveness of our loss, we rely on various tasks: extractive question answering, generative question answering, and summarization. Our evaluation spans multiple widely-used teacher-student pairs, with diverse vocabularies and model architectures.
@@
    Contributing to future research. We make our code [^1] https://github.com/Nicolas-BZRD/llm-recipes [^2] https://github.com/Nicolas-BZRD/llm-distillation, model weights, and generated datasets [^3] https://huggingface.co/Nicolas-BZRD openly available to facilitate future research, minimizing computational overhead and lowering entry barriers.
@@ 2.1 Notations
We define $\Omega$ a model vocabulary set, $|\Omega|$ his size, and $\Omega^{*}$ signifies its Kleene closure [^4] The Kleene closure corresponds to sequences of arbitrary size written with words in $\Omega$. Formally: $\Omega^{*}=\bigcup_{i=0}^{\infty}\Omega^{i}$.. Let $\mathcal{P}(\Omega)$ denote the probability distributions set defined over the sample space $\Omega$. It is defined as $\mathcal{P}(\Omega)=\mathbf{p}\in[0,1]^{\Omega}$ with $\sum_{j=1}^{\Omega}p_{j}=1$. Consider $\mathcal{D}$ as a non-empty set with independent and identically distributed samples ${(\mathbf{x}^{i})}_{i=1}^{N}\in\Omega^{*}$. Each $\mathbf{x}^{i}$ is a sequence of tokens, where $x^{i}_{j}\in\Omega$ represents the $j$th token of the $i$th sequence. The notation $\mathbf{x}^{i}_{<t}=(x^{i}_{t-1},\ldots,x^{i}_{0})\in\Omega^{*}$ denotes the prefix of length $t$. In this paper, we will denote by $\Omega^{S}$ the student vocabulary and by $\Omega^{T}$ the teacher’s vocabulary.
@@ 2.2 Knowledge Distillation Framework
In knowledge distillation (KD), the objective is to guide the learning of a student model using a more complex teacher model Buciluundefined et al. (2006); Hinton et al. (2015). Generally, this paradigm comprises two key components: a cross-entropy loss ($\mathcal{L}_{CE}$), which guides the student model to predict gold tokens accurately, and a secondary loss tasked to align the probability distributions of the student model with those of the teacher model to facilitate the transfer of general knowledge ($\mathcal{L}_{KD}$). Formally, the goal of the student is to minimize $\mathcal{L}$:
@@ 2.3 Knowledge Distillation Related Work
Distillation from teacher-generated text occurs when $\lambda=0$. This strategy is particularly advantageous when dealing with synthetic data Kramchaninova and Defauw (2022); Du et al. (2023); Ushio et al. (2023), a fact highlighted by the effectiveness of instructing large language models such as GPT-3.5/4 Wu et al. (2023); Bubeck et al. (2023). Distillation from teacher-generated text Kim and Rush (2016); He et al. (2023); Hsieh et al. (2023); Zhou and Chiam (2023) will be considered as a baseline throughout the paper. The primary drawback of these methods lies in their failure to fully leverage all the information that can be provided by the teacher.
@@
A further refinement of Eq. 1 occurs when $\mathcal{L}_{KD}$ relies on the internal features of the teacher to transfer knowledgeJiao et al. (2020); Sun et al. (2020). Popular features include transformer attention and internal layers within both encoder-only and encoder-decoder models Raman et al. (2023); Wang et al. (2020a, 2021). A main drawback is that these methods demand access to the models’ internal mechanisms, which is typically unavailable through API access. Moreover, they often assume similarities in architectural patterns between the teacher and the student.
@@
In the black-box approach, practitioners are limited to accessing only the output probabilities or logits of the model. They use these logits to align the student’s output probabilities with those of the teacher through Kullback–Leibler divergence (KL) Sanh et al. (2020). This method has emerged as one of the most widely adopted approaches, successfully distilling encoder, decoder, or encoder-decoder models Timiryasov and Tastet (2023); Mohammadshahi et al. (2022); Zhao et al. (2023a). However, employing KL divergence necessitates that both student and teacher share the same vocabulary, a requirement impractical with current large language models (LLMs) as reported in Fig. 1. We dig into the limitations of this method in the next section.
@@ 2.4 KL Distillation loss
When distilling using the KL Li et al. (2021), the goal is to force the student to learn the teacher’s output probability distribution at each generation step. The formal definition of the objective function is provided in Eq. 2.
@@
Eq. 2 relies on equality across the vocabulary of the student and the teacher, i.e. $\Omega=\Omega^{S}=\Omega^{T}$, enabling us to compute the KL divergence by ensuring similar support of probabilities.
@@
Our examination of Eq. 2 illustrating in Fig. 2 highlights that both vocabulary and absolute continuity constraints pose challenges to distilling two distinct LLM families with logits using KL loss. In the following section, we introduce our ULD loss, providing a flexible framework for knowledge distillation across a wide range of LLMs.
@@ 3.1 Background on Optimal Transport
Optimal transport mathematically transfers mass between distributions, minimizing cost Villani et al. (2009); Peyré et al. (2019). In this context, Wasserstein distance, or Earth Mover Distance, robustly measures dissimilarities between distributions. This distance metric has gained popularity in NLP applications such as hallucination detection Guerreiro et al. (2022); Shuster et al. (2021), clustering Zhuang et al. (2022); Ye et al. (2017); Xu et al. (2018) or sentence similarity:Colombo et al. (2021); Xu et al. (2018); Bahuleyan et al. (2018).
@@ 3.2 Universal Logit Distillation loss
The ULD loss retains the CE loss term to guide the model in generating the target token and introduces a Wasserstein Distance term to transfer knowledge from the teacher to the student. By minimizing the distance between the soft probabilities of the teacher and the student, our goal is to reproduce not only the predictions for the golden token but also the near-zero labels, which are crucial for performance and generalization.
@@
where $\mathbf{|x|}=min\left(|\mathbf{x}^{S}|,|\mathbf{x}^{T}|\right)$ the minimum length between the sentences tokenized with the teacher or student tokenizers. Respectively $\mathbf{q_{\bm{\theta}_{T}}}\left(\cdot|\mathbf{x}^{T}_{<t}\right)$ and $\mathbf{p_{\bm{\theta}_{S}}}\left(\cdot|\mathbf{x}^{S}_{<t}\right)$ are the probability distribution for the Teacher and Student models at the $t^{th}$ steps, $\text{CE}(t)$ the cross-entropy loss defined in Sec. 2.4 and $\mathcal{W}_{1}$ represents the discrete Wasserstein distance defined in Eq. 3 where $\lambda\in\mathbb{R}^{+}$ controls the trade-off between the two terms and set to $1.5$ in the rest of this paper.
@@ 3.3 Fast Computation & Approximations
To the best of our knowledge, we are the first to motivate and propose the Wasserstein distance as a learning loss for distillation in the scope of the LLM decoder. Prior efforts focus on Sinkhorn regularization Cuturi (2013) with encoder-decoder for classification Bhardwaj et al. (2022), while our focus diverges as we concentrate on the generative setting. This shift presents inherent challenges, as the naive computation of the Wasserstein loss in Eq. 3 exhibits a complexity of $\mathcal{O}(n^{3}\log n)$, where $n$ signifies the size of the larger support. While manageable in small classification scenarios with encoders, the magnitude of the vocabulary, which can extend to 100K tokens in generative tasks, renders this approach intractable, particularly for long sequences.
@@
Uniform Support Length: We augment either the student or teacher vocabulary size through distribution padding (with 0 value), ensuring equal support size for both (i.e., $|\Omega_{t}|=|\Omega_{s}|=|\Omega|$).
@@
Under this assumption the Wasserstein distance used in the $\mathcal{L}_{\text{ULD}}$ loss becomes Peyré et al. (2019):
@@
where $\sigma^{S}$ and $\sigma^{T}$ as the permutation that sorts the decreasing probability of student and teacher soft probability vectors.
@@
The final version of Eq. 5 is straightforward: with each generation, we pad and sort the probability vectors of both student and teacher, then compute the absolute difference. This ensures that no matter the token, the confidence levels of the sorted teacher and student are close.

### v3

@@ Abstract
Deploying large language models (LLMs) with billions of parameters is often impractical in industrial settings due to constraints like cost, latency, and hardware limitations. Knowledge distillation (KD) provides a solution by compressing the knowledge from large, resource-intensive models into task-specific smaller ones. Various strategies exist, some relying on the text generated by the teacher model, optionally, leveraging its output logits to improve learning. However, these logit-based methods usually require the teacher and student models to share the same tokenizer, which limits their applicability across different model families. In this paper, we propose the Universal Logit Distillation (ULD) loss, which uses optimal transport theory to enable distillation across different architectures and tokenizers. Our results demonstrate that ULD loss effectively facilitates the distillation process, paving the way for a more widespread use of distillation.
@@ 1 Introduction
A significant trend in NLP involves the utilization of large language models (LLMs) such as LLama (Touvron et al., 2023a), Mistral (Jiang et al., 2023), Falcon (Almazrouei et al., 2023), GPT-NeoX (Black et al., 2022), or Mixtral (Jiang et al., 2024). While LLMs offer impressive performance (Bubeck et al., 2023), their deployment is often hampered by hardware availability, cost, and latency bottlenecks. Several strategies, such as efficient decoding (Leviathan et al., 2023; Ye et al., 2023), model recycling (Lester et al., 2022), and model size reduction (Dettmers et al., 2023; Ma et al., 2023), have been developed to streamline their use. Among these, knowledge distillation (KD) (Buciluundefined et al., 2006; Hinton et al., 2015) a widely adopted technique (Sanh et al., 2020; Jiao et al., 2020; Mohammadshahi et al., 2022; He et al., 2023; Raman et al., 2023; Dasgupta et al., 2023), transferring the capabilities of large, complex teacher models into more manageable and smaller student models, tailored for specific tasks. This approach is particularly valuable in contexts where the comprehensive knowledge contained within LLMs is not wholly necessary. This process aims to maintain the peak performance of general models on specific tasks while minimizing latency and memory usage.
Figure 1: Vocabulary overlap between teacher and student models (Sec. 4.2), illustrating the challenge of cross-tokenizer distillation. (e.g., Bloomz tokenizer has 30.03% of Mistral’s vocabulary).
Two approaches can be considered. The "white box" approach, where researchers propose loss functions that require access to the model architecture to compute similarities across layers, forcing adjustment for each situation. In contrast, the "black box" approach, indifferent to models latent spaces, relies solely on the output logit vectors from teacher and student. The black-box approach, due to its flexibility and generality, is easily implemented by practitioners through libraries or APIs, facilitating its adoption.
Over the past years, NLP researchers have extensively explored these distillation methods with teacher and student models sharing similar architecture, notably the BERT encoder (Sanh et al., 2020; Jiao et al., 2020; Sun et al., 2020). This desire to maintain similar architecture between teacher and student models by mirroring some of the teacher blocks, hidden sizes, or relying on the same tokenizer, came from the need to find similar information supports to apply distillation losses.
@@
However, KD for generative models, those relying on encoder-decoder or decoder architectures, has received less attention due to the lack of common support between models. Although models of different sizes exist, they often diverge in architecture and tokenizer (Fig. 1), making logit distillation loss inapplicable. In fact, recent research predominantly focuses on synthetic data fine-tuning (He et al., 2022; Kramchaninova & Defauw, 2022; Ouyang et al., 2022), rather than refining logits loss in the black box approach. Thus far, the primary method for KD with decoder models is to use the text generated by the teacher model (He et al., 2023; Hsieh et al., 2023), and if possible, when students and teachers models belong to the same family, improving KD by employing output-logit distillation with Kullback–Leibler divergence (Mohammadshahi et al., 2022; Raman et al., 2023; Wang et al., 2020a). This raises the following research question:
How can we build a general distillation loss that leverages logits capacity while staying within a black box framework that is easy to implement?
@@
    A universal logit distillation loss. We introduce a new loss, Universal Logit Distillation Loss (ULD loss), versatile to tokenizers and with minimal assumptions about the architectures of teacher and student models.
@@
    Experimental Results. We demonstrate the robust effectiveness of our loss function in transferring the capabilities of many teacher models to different smaller student models for a variety of specific tasks, including extractive question answering, generative question answering, and summarization.
@@
    Contributing to future research. We make our code [^1] https://github.com/Nicolas-BZRD/llm-recipes, model weights, and generated datasets [^2] https://huggingface.co/Nicolas-BZRD openly available to facilitate future research, minimizing computational overhead and lowering entry barriers.
@@ 2.1 Notations
We define $\Omega$ a model vocabulary set, $|\Omega|$ his size, and $\Omega^{*}$ signifies its Kleene closure [^3] The Kleene closure corresponds to sequences of arbitrary size written with words in $\Omega$. Formally: $\Omega^{*}=\bigcup_{i=0}^{\infty}\Omega^{i}$.. Let $\mathcal{P}(\Omega)$ denote the set of probability distributions defined over the sample space $\Omega$. It is defined as $\mathcal{P}(\Omega)=\mathbf{p}\in[0,1]^{\Omega}$ with $\sum_{j=1}^{\Omega}p_{j}=1$. Consider $\mathcal{D}$ as a non-empty set with independent and identically distributed samples ${(\mathbf{x}^{i})}_{i=1}^{N}\in\Omega^{*}$. Each $\mathbf{x}^{i}$ is a sequence of tokens, where $x^{i}_{j}\in\Omega$ represents the $j$th token of the $i$th sequence. The notation $\mathbf{x}^{i}_{<t}=(x^{i}_{t-1},\ldots,x^{i}_{0})\in\Omega^{*}$ denotes the prefix of length $t$. In this paper, we will denote by $\Omega^{S}$ the student vocabulary and by $\Omega^{T}$ the teacher’s vocabulary.
@@ 2.2 Knowledge Distillation Framework
In knowledge distillation (KD), the objective is to guide the learning of a student model on a specific task using a complex teacher model (Buciluundefined et al., 2006; Hinton et al., 2015) that possesses generic knowledge. This paradigm includes two main components: a cross-entropy loss ($\mathcal{L}{CE}$), which ensures the student model accurately predicts the gold tokens, and a secondary loss ($\mathcal{L}{KD}$) tasked to align the probability distributions of the student model with those of the teacher model. This alignment transfers broader knowledge not encapsulated in gold vectors (similar to standard basis vectors) used in the cross-entropy loss. Formally, the goal of the student is to minimize $\mathcal{L}$:
@@ 2.3 Knowledge Distillation Related Work
Distillation from teacher-generated text occurs when $\lambda=0$. This strategy is particularly advantageous when dealing with synthetic data (Kramchaninova & Defauw, 2022; Du et al., 2023; Ushio et al., 2023), a fact highlighted by the effectiveness of instructing large language models such as GPT-3.5/4 (Wu et al., 2023; Bubeck et al., 2023). Distillation from teacher-generated text (Kim & Rush, 2016; He et al., 2023; Hsieh et al., 2023; Zhou & Chiam, 2023) will be considered as a baseline throughout the paper. The primary drawback of these methods lies in their failure to fully leverage all the information that can be provided by the teacher.
@@
A further refinement of Eq. 1 occurs when $\mathcal{L}_{KD}$ relies on the internal features of the teacher to transfer knowledge(Jiao et al., 2020; Sun et al., 2020). Popular features include transformer attention and internal layers within both encoder-only and encoder-decoder models (Raman et al., 2023; Wang et al., 2020a; 2021). However, these methods require access to models’ internal mechanisms, which are not available through API access, and assume similarities in architectural patterns between models, forcing adjustments for each situation.
@@
In the black-box approach, practitioners only used the output logits of the model. They use these logits to align the student’s output probabilities with those of the teacher through Kullback–Leibler divergence (KL) (Sanh et al., 2020). This method has emerged as one of the most widely adopted approaches, successfully distilling encoder, decoder, or encoder-decoder models (Timiryasov & Tastet, 2023; Mohammadshahi et al., 2022; Zhao et al., 2023a). However, employing KL divergence necessitates that both student and teacher share the same vocabulary, a requirement impractical with current large language models (LLMs) as reported in Fig. 1. We dig into the limitations of this method in the next section.
@@ 2.4 KL Distillation loss
When distilling using the KL (Li et al., 2021), the goal is to force the student to learn the teacher’s output probability distribution at each generation step. The formal definition of the objective function is provided in Eq. 2.
@@
Eq. 2 relies on equality across the vocabulary of the student and the teacher, i.e. $\Omega=\Omega^{S}=\Omega^{T}$, ensuring similar support of probability distributions.
@@
Our examination of Eq. 2 illustrating in Fig. 2 highlights that both vocabulary (similar support of probability distributions) and absolute continuity constraints pose challenges to distilling two distinct LLM families with logits using KL loss. In the following section, we introduce our ULD loss, providing a flexible framework for knowledge distillation across a wide range of LLMs.
@@ 3.1 Background on Optimal Transport
Optimal transport provides a mathematical framework for transferring mass between distributions while minimizing cost (Villani et al., 2009; Peyré et al., 2019). Within this framework, the Wasserstein distance, also known as the Earth Mover Distance, is a robust metric for quantifying dissimilarities between distributions. This distance metric has gained popularity in NLP applications such as hallucination detection (Guerreiro et al., 2022; Shuster et al., 2021), clustering (Zhuang et al., 2022; Ye et al., 2017; Xu et al., 2018) or sentence similarity (Colombo et al., 2021; Xu et al., 2018; Bahuleyan et al., 2018).
@@ 3.2 Universal Logit Distillation loss
The ULD loss retains the CE loss term to guide the model in generating the target token and introduces a Wasserstein Distance term to transfer knowledge from the teacher to the student. By minimizing the distance between the soft probabilities of the teacher and the student, our goal is to reproduce not only the predictions for the gold token but also the near-zero labels, which are crucial for performance and generalization.
@@
where $\mathbf{|x|}=min\left(|\mathbf{x}^{S}|,|\mathbf{x}^{T}|\right)$ the minimum length between the sentences tokenized with the teacher or student tokenizers. Respectively $\mathbf{q_{\bm{\theta}_{T}}}\left(\cdot|\mathbf{x}^{T}_{<t}\right)$ and $\mathbf{p_{\bm{\theta}_{S}}}\left(\cdot|\mathbf{x}^{S}_{<t}\right)$ are the probability distribution for the Teacher and Student models at the $t^{th}$ steps, $\text{CE}(t)$ the cross-entropy loss defined in Sec. 2.4 and $\mathcal{W}_{1}$ represents the discrete Wasserstein distance defined in Eq. 3 where $\lambda\in\mathbb{R}^{+}$ controls the trade-off between the two terms and set to $1.5$ in the rest of this paper as discussed in Appendix: Sec. B.1.
@@ 3.3 Fast Computation & Approximations
To the best of our knowledge, we are the first to motivate and propose the Wasserstein distance as a learning loss for distillation in the scope of the LLM decoder. Prior efforts focus on Sinkhorn regularization (Cuturi, 2013) with encoder-decoder for classification (Bhardwaj et al., 2022), while our focus diverges as we concentrate on the generative setting. This shift presents inherent challenges, as the naive computation of the Wasserstein loss in Eq. 3 exhibits a complexity of $\mathcal{O}(n^{3}\log n)$, where $n$ signifies the size of the larger support. While manageable in small classification scenarios with encoders, the magnitude of the vocabulary, which can extend to 100K tokens in generative tasks, renders this approach intractable, particularly for long sequences.
@@
Uniform Support Length: We augment either the student or teacher vocabulary size through distribution padding, ensuring equal support size for both (i.e., $|\Omega_{t}|=|\Omega_{s}|=|\Omega|$).
@@
Under this assumption the Wasserstein distance used in the $\mathcal{L}_{\text{ULD}}$ loss becomes the one introduced by Peyré et al. (2019):
@@
with a complexity of $\mathcal{O}\left(n\cdot\log(n)\right)$ (Sec. A.1) where $\sigma^{S}$ and $\sigma^{T}$ are the permutations that sort in decreasing order the probability of student and teacher probability vectors. An extensive demonstration of this closed form is provided in App. A, written clearly and accessible.
@@
The final version of Eq. 5 is decomposed into two sums:
  • The outer sum iterates over the sequence length (denoted as $|\mathbf{x}|$) to compute the Wasserstein distance at each time step $t$.
  • The inner sum $\sum_{i}|p_{\sigma^{S}(s_{i})}-q_{\sigma^{T}(t_{i})}|$ represents the Wasserstein ($\mathcal{W}_{1}$) distance between the logits of the student and teacher models at that time step.
To compute the $\mathcal{W}_{1}$ distance efficiently, we use a closed-form solution derived under certain assumptions: uniform support length and uniform transport cost, as detailed in App. A. This solution computes the absolute difference between the sorted probability masses of the teacher and student vectors at time step $t$, ensuring that the overall structure of the teacher’s probability distribution is preserved.
This process allows the student model to learn not only from the golden tokens (as used in the Cross-Entropy Loss) but also from the surrounding probability distribution, capturing distributional information from the teacher’s predictions.